In the month of March, Style Salon services 570 clients at an average price of $170. During the month, fixed costs were $29,784 and variable costs were 60% of sales.

Determine the contribution margin in dollars, per unit, and as a ratio. (Round answers to 0 decimal places, e.g. 1,225.)
Contribution margin $

Contribution margin per unit $

Contribution margin ratio

%
Using the contribution margin technique, compute the break-even point in dollars and in units. (Round answers to 0 decimal places, e.g. 1,225.)
Break-even sales $

Break-even sales

units
Answer #1

Contribution margin

= Sales - Variable costs

= (570*170)-60%

38,760

Contribution margin per unit

= (38,760/570)

68

Contribution margin ratio

= (68/170)

40%

.

Breakeven sales = Fixed cost/Contribution margin ratio

= 29,784/40%

74,460

Breakeven sales = Fixed cost/Contribution margin per unit

= 29,784/68

438

Comment if you face any issues